# Tight Algorithm for Connected Odd Cycle Transversal Parameterized by Clique-width

### 12 Feb 2024 v1

@@ 5.3 The Algorithm counts action-sequences
In this section, we show that the tables $T$ defined in Section 4 count (modulo 2) the number of action-sequences of weight ${w}$, that generate patterns consistent with a complete pattern $q\in\mathscr{P}_{C}$.
@@
We say that a set of patterns $S$ parity-represents another set of patterns $T$ over a family of patterns $\mathscr{P}^{*}\subseteq\mathscr{P}$, if and only if for each pattern $q\in\mathscr{P}^{*}$ it holds that
@@
We say that $S$ parity-represents a pattern $p$ over $\mathscr{P}^{*}$, when $S$ parity-represents $\{p\}$ over $\mathscr{P}^{*}$.
@@
The operations defined in Definition 3.2 preserve $\mathscr{P}_{C}$-parity-representation over $\mathscr{P}_{C}$.
@@
Let $S,T\subseteq\mathscr{P}_{C}$ with $T$ parity-represents $S$ over $\mathscr{P}_{C}$, and $i,j\in[k]$. Then it holds for all $\ell\in[4]$ that $\overset{\triangle}{\operatorname{Ac}}(\square_{i,j}T,\ell)$ parity-represents $\overset{\triangle}{\operatorname{Ac}}(\square_{i,j}S,\ell)$ over $\mathscr{P}_{C}$.
@@
Let $\operatorname{Op}$ be a $k$-ary operation that preserves $\mathscr{P}^{*}$-parity-representation over $\hat{\mathscr{P}}\subseteq\mathscr{P}$ for two families $\mathscr{P}^{*},\hat{\mathscr{P}}$. Let $P_{1},P_{2}\subseteq\hat{\mathscr{P}}$ be two subfamilies, both closed under $\operatorname{Op}$. Let $S_{1},\dots S_{k}\in\{0,1\}^{P_{1}}$, and $T_{1},\dots T_{k}\in\{0,1\}^{P_{2}}$. If it holds for each $i\in[k]$, and for each $q\in\mathscr{P}^{*}$ that
@@
Hence, it holds that $X_{i}$ parity-represents $Y_{i}$ over $\mathscr{P}^{*}$ for each $i\in[k]$. Let $X^{*}=\overset{\triangle}{\operatorname{Op}}(X_{1},\dots,X_{k})$, and $Y^{*}=\overset{\triangle}{\operatorname{Op}}(Y_{1},\dots,Y_{k})$. Since $\operatorname{Op}$ preserves $\mathscr{P}^{*}$-parity representation over $\hat{\mathscr{P}}$, it holds that $X^{*}$ parity-represents $Y^{*}$ over $\mathscr{P}^{*}$. We claim that
@@
For each node $x\in V(\mathcal{T})$, and all values ${b},{w}$, and all states $c\in\mathscr{C}$ and $q\in\mathscr{P}_{C}$, it holds that
@@
if and only if there exist odd many valid action-sequences $\tau$ at $x$ of cost ${b}$ and weight ${w}$ generating the pair $(p,c)$, for some complete pattern $p\in\mathscr{P}_{C}$ with $p\sim q$.
@@
For $p\in\mathscr{P}_{C}$ and $c\in\mathscr{C}$, let $D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}$ be the number of valid action-sequences at $x$ of cost ${b}$ and weight ${w}$ generating the pair $(p,c)$. We prove by induction over $\mathcal{T}$ that for all $x\in\mathcal{V}$, all values ${b},{w}$, for all $c\in C$ and all complete patterns $q\in\mathscr{P}_{C}$ it holds that
@@
  • Create node $\mu_{x}=i(v)$: Clearly there exist four valid action-sequences at $x$, each generating a complete pattern that is a $CS$-pattern as well. Hence, it holds for $p\in\mathscr{P}_{CS}$ that
@@
  • Relabel node $\mu_{x}=\rho_{i\rightarrow j}(\mu_{x^{\prime}})$: Let $(p,c)\in\mathscr{P}_{C}\times\mathscr{C}$. For each pair $(p^{\prime},c^{\prime})\in\mathscr{P}_{C}\times\mathscr{C}$ with $(p^{\prime})_{i\rightarrow j}=p$ and $(c^{\prime})_{i\rightarrow j}=c$ it holds that each valid action-sequence at $x^{\prime}$ generating $(p^{\prime},c^{\prime})$, is a valid action-sequence at $x$ that generates $(p,c)$. Also, for each valid action-sequence $\tau$ at $x$ generating $(p,c)$, there exists such $p^{\prime},c^{\prime}$ with the aforementioned property, such that $\tau_{x^{\prime}}$ generates $(p^{\prime},c^{\prime})$ at $x^{\prime}$. Hence, it holds that
@@
    By induction hypothesis, it holds for all ${b},{w}$, $c\in\mathscr{C}$ and $q\in\mathscr{P}_{C}$ that
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
  • Union node $\mu_{x}=\mu_{x_{1}}\uplus\mu_{x_{2}}$: Let $(p,c)\in\mathscr{P}_{C}\times\mathscr{C}$. For all pairs $(p_{1},c_{1}),(p_{2},c_{2})\in\mathscr{P}_{C}\times\mathscr{C}$ with $p_{1}\oplus p_{2}=p$ and $c_{1}\oplus_{C}c_{2}=c$ it holds that for each two valid action-sequences $\tau_{1}$ generating $(p_{1},c_{1})$ at $x_{1}$, and $\tau_{2}$ generating $(p_{2},c_{2})$ at $x_{2}$, that $\tau=\tau_{1}\dot{\cup}\tau_{2}$ is a valid action-sequence at $x$ that generates $(p,c)$. Also, for each valid action-sequence $\tau$ at $x$ generating $(p,c)$, there exists such $p_{1},p_{2},c_{1},c_{2}$ with $p_{1}\oplus p_{2}=p$ and $c_{1}\oplus_{C}c_{2}=c$, such that $\tau_{x_{1}}$ generates $(p_{1},c_{1})$ at $x_{1}$ and $\tau_{x_{2}}$ generates $(p_{2},c_{2})$ at $x_{2}$. Hence, it holds that
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
  • Join node $\mu_{x}=\eta_{i,j}(\mu_{x^{\prime}})$: For $c\in\mathscr{C}$ with $c(i)\not\sim c(j)$, it holds clearly for each complete pattern $p$, that
    $D{\big{[}{x,{b},{w}}\big{]}}\big{[}{p,c}\big{]}=0.$
    Since the same holds for $T$ and each $\mathscr{P}_{CS}$ pattern by definition, the claim follows in this case. Now assume that $c(i)\sim c(j)$. For $\ell\in[4]$, let $\tau$ be a valid action-sequence at $x$ generating the pair $(p,c)$ for some $p\in\mathscr{P}_{C}$, such that $\tau(x)=\ell$. It must holds that $\tau_{x^{\prime}}$ is a valid action-sequence at $x^{\prime}$ generating the pair $(p^{\prime},c)$ with $\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)=p$. Moreover, since we assume that $c(i)\sim c(j)$, each valid action-sequence $\tau^{\prime}$ at $x^{\prime}$ generating the pair $(p^{\prime},c)$ for some $p^{\prime}\in\mathscr{P}_{C}$ can be extended by $x\mapsto\ell$ to a valid action-sequence at $x$ generating the pair $(p,c)$ with $p=\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)$. Hence, it holds that
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
    We define the tables $T^{\prime}{\big{[}{x^{\prime},{b},{w}}\big{]}}\in\{0,1\}^{\mathscr{P}_{C}}$ as
@@
    Then clearly, it holds for all $q\in\mathscr{P}_{C}$ that
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}T^{\prime}{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
    Finally, let $T^{\prime\prime}{\big{[}{x,{b},{w}}\big{]}}\in\{0,1\}^{\mathscr{P}_{C}}$ defined by
@@
    Then it holds by Lemma 5.19 and Lemma 5.18 that
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}T^{\prime\prime}{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}.$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}T^{\prime\prime}{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(c,p)}\big{]},$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}\sum\limits_{\ell\in[4]}\sum\limits_{\begin{subarray}{c}p^{\prime}\in\mathscr{P}_{CS}\\ p\in\operatorname*{PREP}\big{(}\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)\big{)}\end{subarray}}T\big{[}{x^{\prime},{b},{w}-\omega(x,\ell)}\big{]}\big{[}{(p^{\prime},c)}\big{]}$
    However, this holds if for each $\ell\in[4]$ it holds that
@@
    which is the case, if for each $p^{\prime}\in\mathscr{P}_{CS}$ and $p=\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)$ it holds that $p\sim q$ if and only if |$|\{r\in\operatorname*{PREP}(p)\colon p\sim q\}|$ is odd, which is clearly the case from Lemma 5.15.
@@
Let ${b},{w}$ be two values, and $c\in\mathscr{C}$. It holds that $T{\big{[}{r,{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=1$ if and only if there exist odd many valid action-sequences of cost ${b}$ and weight ${w}$ at $r$ generating $([0],c)$.
@@
$\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim[0]\end{subarray}}T{\big{[}{r,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}=1,$
if and only if there exist odd many valid action-sequences $\tau$ at $r$ of cost ${b}$ and weight ${w}$ generating a pair $(p,c)$ with $p\sim[0]$. It holds by Lemma 3.6 that $[0]$ is the only complete pattern consistent with $[0]$, and hence, it holds that
$T{\big{[}{r,{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim[0]\end{subarray}}T{\big{[}{r,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}=1,$
if and only if there are exist odd many valid action-sequences $\tau$ at $r$ of cost ${b}$ and weight ${w}$ generating a pair $(p,c)$ with $p\sim[0]$. ∎

### 26 Feb 2024 v2

@@ 5.3 The Algorithm counts action-sequences
In this section, we show that using the tables $T$ defined in Section 4, one can count (modulo 2) the number of action-sequences of cost ${b}$ and weight ${w}$, that generate patterns consistent with a complete pattern $q\in\mathscr{P}_{C}$.
@@
We say that a set of patterns $S$ parity-represents another set of patterns $T$ over a family of patterns $\mathscr{P}^{*}\subseteq\mathscr{P}$, if for each pattern $q\in\mathscr{P}^{*}$ it holds that
@@
We say that $S$ parity-represents a pattern $p$ over $\mathscr{P}^{*}$, if $S$ parity-represents $\{p\}$ over $\mathscr{P}^{*}$.
@@
The operations join, relabel and union, as defined in Definition 3.2, preserve $\mathscr{P}_{C}$-parity-representation over $\mathscr{P}_{C}$.
@@
Let $S,T\subseteq\mathscr{P}_{C}$ be two families of complete patterns, such that $T$ parity-represents $S$ over $\mathscr{P}_{C}$, and $i,j\in[k]$. Then it holds for all $\ell\in[4]$ that $\overset{\triangle}{\operatorname{Ac}}(\square_{i,j}T,\ell)$ parity-represents $\overset{\triangle}{\operatorname{Ac}}(\square_{i,j}S,\ell)$ over $\mathscr{P}_{C}$.
@@
Let $\operatorname{Op}$ be some $k$-ary operation over patterns that preserves $\mathscr{P}^{*}$-parity-representation over $\hat{\mathscr{P}}\subseteq\mathscr{P}$ for two families $\mathscr{P}^{*},\hat{\mathscr{P}}$. Let $P_{1},P_{2}\subseteq\hat{\mathscr{P}}$ be two subfamilies, both closed under $\operatorname{Op}$. Let $S_{1},\dots S_{k}\in\{0,1\}^{P_{1}}$, and $T_{1},\dots T_{k}\in\{0,1\}^{P_{2}}$. If it holds for each $i\in[k]$, and for each $q\in\mathscr{P}^{*}$ that
@@
Hence, it holds that $X_{i}$ parity-represents $Y_{i}$ over $\mathscr{P}^{*}$ for each $i\in[k]$. Let $X^{*}=\overset{\triangle}{\operatorname{Op}}(X_{1},\dots,X_{k})$, and $Y^{*}=\overset{\triangle}{\operatorname{Op}}(Y_{1},\dots,Y_{k})$. Since $\operatorname{Op}$ preserves $\mathscr{P}^{*}$-parity-representation over $\hat{\mathscr{P}}$, it holds that $X^{*}$ parity-represents $Y^{*}$ over $\mathscr{P}^{*}$. We claim that
@@
For each node $x\in V(\mathcal{T})$, for all values ${b},{w}$, and all $c\in\mathscr{C}$ and $q\in\mathscr{P}_{C}$, it holds that
@@
if and only if there exists an odd number of valid action-sequences $\tau$ at $x$ of cost ${b}$ and weight ${w}$ generating the pair $(p,c)$, where $p\sim q$.
@@
For $p\in\mathscr{P}_{C}$ and $c\in\mathscr{C}$, let $D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}$ be the number of valid action-sequences at $x$ of cost ${b}$ and weight ${w}$ generating the pair $(p,c)$. We prove by induction over $\mathcal{T}$ that for all $x\in\mathcal{V}$, for all values ${b},{w}$, for all $c\in\mathscr{C}$ and all complete patterns $q\in\mathscr{P}_{C}$ it holds that
@@
  • Introduce node $\mu_{x}=i(v)$: Clearly there exist four valid action-sequences at $x$, each generating a complete pattern that is a $CS$-pattern as well. Hence, it holds for $p\in\mathscr{P}_{CS}$ that
@@
  • Relabel node $\mu_{x}=\rho_{i\rightarrow j}(\mu_{x^{\prime}})$: Let $(p,c)\in\mathscr{P}_{C}\times\mathscr{C}$. For each pair $(p^{\prime},c^{\prime})\in\mathscr{P}_{C}\times\mathscr{C}$ with $(p^{\prime})_{i\rightarrow j}=p$ and $(c^{\prime})_{i\rightarrow j}=c$ it holds that each valid action-sequence at $x^{\prime}$ generating $(p^{\prime},c^{\prime})$, is a valid action-sequence at $x$ that generates $(p,c)$. It also holds for each valid action-sequence $\tau$ at $x$ generating $(p,c)$, that $\tau_{x^{\prime}}$ generates a pair $(p^{\prime},c^{\prime})$ at $x^{\prime}$ where $(p^{\prime})_{i\rightarrow j}=p$ and $(c^{\prime})_{i\rightarrow j}=c$. Hence, it holds that
@@
    By induction hypothesis, it holds for all values ${b},{w}$, and all $c\in\mathscr{C}$ and $q\in\mathscr{P}_{C}$ that
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
  • Union node $\mu_{x}=\mu_{x_{1}}\uplus\mu_{x_{2}}$: Let $(p,c)\in\mathscr{P}_{C}\times\mathscr{C}$. For all pairs $(p_{1},c_{1}),(p_{2},c_{2})\in\mathscr{P}_{C}\times\mathscr{C}$ with $p_{1}\oplus p_{2}=p$ and $c_{1}\oplus_{C}c_{2}=c$ it holds that for each two valid action-sequences $\tau_{1}$ generating $(p_{1},c_{1})$ at $x_{1}$, and $\tau_{2}$ generating $(p_{2},c_{2})$ at $x_{2}$, that $\tau=\tau_{1}\dot{\cup}\tau_{2}$ is a valid action-sequence at $x$ that generates $(p,c)$. Also, for each valid action-sequence $\tau$ at $x$ generating $(p,c)$, there exist two tuples $(p_{1},c_{1}),(p_{2},c_{2})$ with $p_{1}\oplus p_{2}=p$ and $c_{1}\oplus_{C}c_{2}=c$, such that $\tau_{x_{1}}$ generates $(p_{1},c_{1})$ at $x_{1}$ and $\tau_{x_{2}}$ generates $(p_{2},c_{2})$ at $x_{2}$. Hence, it holds that
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
  • Join node $\mu_{x}=\eta_{i,j}(\mu_{x^{\prime}})$: For $c\in\mathscr{C}$ with $c(i)\not\sim c(j)$, it holds clearly for each complete pattern $p$, that $D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}=0$. Since it also holds that $T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}=0$ for each $p\in\mathscr{P}_{CS}$ by definition, the claim follows in this case. Now assume that $c(i)\sim c(j)$. For $\ell\in[4]$, let $\tau$ be a valid action-sequence at $x$ generating the pair $(p,c)$ for some $p\in\mathscr{P}_{C}$, such that $\tau(x)=\ell$. It must hold that $\tau_{x^{\prime}}$ is a valid action-sequence at $x^{\prime}$ generating the pair $(p^{\prime},c)$ with $\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)=p$. Moreover, since we assume that $c(i)\sim c(j)$, each valid action-sequence $\tau^{\prime}$ at $x^{\prime}$ generating the pair $(p^{\prime},c)$ for some $p^{\prime}\in\mathscr{P}_{C}$ can be extended by $x\mapsto\ell$ to a valid action-sequence at $x$ generating the pair $(p,c)$ with $p=\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)$. Hence, it holds that
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
    We define the tables $T^{\prime}{\big{[}{x^{\prime},{b},{w}}\big{]}}\in\{0,1\}^{\mathscr{P}_{C}\times\mathscr{C}}$ as
@@
    Then clearly, it holds for all values ${b},{w}$, and all $q\in\mathscr{P}_{C}$ that
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}T^{\prime}{\big{[}{x^{\prime},{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
    Finally, let $T^{\prime\prime}{\big{[}{x,{b},{w}}\big{]}}\in\{0,1\}^{\mathscr{P}_{C}\times\mathscr{C}}$ defined by
@@
    It holds by Lemma 5.19 that actions preserve $\mathscr{P}_{C}$-parity-representation over $\mathscr{P}_{C}$, and hence, by Lemma 5.20 that
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}D{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}T^{\prime\prime}{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}.$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{C}\\ p\sim q\end{subarray}}T^{\prime\prime}{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}T{\big{[}{x,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]},$
@@
    $\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim q\end{subarray}}\sum\limits_{\ell\in[4]}\sum\limits_{\begin{subarray}{c}p^{\prime}\in\mathscr{P}_{CS}\\ p\in\operatorname*{PREP}\big{(}\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)\big{)}\end{subarray}}T\big{[}{x^{\prime},{b},{w}-\omega(x,\ell)}\big{]}\big{[}{(p^{\prime},c)}\big{]}.$
    This holds, if for each $\ell\in[4]$ it holds that
@@
    which is the case, if for each $p^{\prime}\in\mathscr{P}_{CS}$ and $p=\operatorname{Ac}(\square_{i,j}p^{\prime},\ell)$ it holds that $p\sim q$ if and only if |$|\{r\in\operatorname*{PREP}(p)\colon p\sim q\}|$ is odd, which is clearly the case due to Lemma 5.16.
@@
For two values ${b},{w}$, and a coloring $c\in\mathscr{C}$, it holds that $T{\big{[}{r,{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=1$ if and only if there exists an odd number of valid action-sequences of cost ${b}$ and weight ${w}$ at $r$ generating $([0],c)$.
@@
$\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim[0]\end{subarray}}T{\big{[}{r,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}1,$
if and only if there exists an odd number of valid action-sequences $\tau$ at $r$ of cost ${b}$ and weight ${w}$ generating a pair $(p,c)$ with $p\sim[0]$. It holds by Lemma 3.6 that $[0]$ is the only complete pattern consistent with $[0]$, and hence, it holds that
$T{\big{[}{r,{b},{w}}\big{]}}\big{[}{([0],c)}\big{]}=\sum\limits_{\begin{subarray}{c}p\in\mathscr{P}_{CS}\\ p\sim[0]\end{subarray}}T{\big{[}{r,{b},{w}}\big{]}}\big{[}{(p,c)}\big{]}\equiv_{2}1,$
if and only if there exists an odd number of valid action-sequences $\tau$ at $r$ of cost ${b}$ and weight ${w}$ generating the pair $([0],c)$. ∎